# The origin of supermassive black holes at cosmic dawn

### v2

@@ 3 Accretion onto a seed black hole in a NFW dark matter halo
In a Hernquist halo, the mass remains finite when integrating the density profile upto infinity. However, for the NFW halo, the mass diverges when integrating the density profile to infinity. Therefore, the mass of the halo is defined upto a virial radius at which the halo profile is truncated. The virial radius is defined as (e.g. Mo et al., 1998),
@@
Locations of the critical points for specific values of $\alpha$ and $\beta$ are shown in Fig. 2 bottom panel.
@@
The outermost critical point is difficult to retrieve analytically. However, we can use the fact that it is approximately at the virial radius or outside it (Fig. 2, see also Hobbs et al. 2012), and the NFW halo is truncated at the virial radius, therefore, the first term on the right hand side of Eq. 23 is simply $1/x_{c}^{2}$, that yields $x_{c}\approx(1+\beta)/2$, same as in case of the Hernquist halo.
In Fig. 3, we show a comparison of the trans-sonic solutions for a Hernquist halo and an NFW halo (for $c=5$). The two types of trans-sonic solutions are shown, one through the innermost critical point (bottom panel), another through the outermost critical point (top panel). Solutions through the innermost point for the two types of dark matter haloes are almost identical because the innermost point is due to the black hole. Moreover the solutions through the outermost critical point also differ only a little between the two haloes.
@@
In general there are three critical points, innermost, intermediate and outermost (see Fig. 2). The Mach number, $\mathcal{M}$, for the accretion flow through the outermost critical point (top panel of Fig. 3), stops increasing as $x$ is closer to the location of the intermediate critical point. Then there are three possibilities depending on the value of $\alpha$. First, for very low $\alpha$, $\mathcal{M}$ increase again as $x$ approaches the location of the innermost critical point. Second, when $\alpha$ is higher and equal to a transition value ($\alpha_{\rm T}$), $\mathcal{M}$ decreases and passes through the innermost critical point. Third, for very high $\alpha$, $\mathcal{M}$ decreases rapidly and the solution cannot reach the innermost critical point, instead it circles around the intermediate critical point.
In the third case, the solution through the outermost critical point cannot reach the black hole. However, the accretion still occurs as the trans-sonic solution through the innermost critical point takes-over (Fig. 3 bottom panel).
The value of $\alpha$ decreases with redshift, therefore, we can understand the switch-over between the two solutions from the perspective of a decreasing $\alpha$ as well. For a very high value of $\alpha$, the feasible trans-sonic solution for the growth of the black hole is through the innermost critical point. As $\alpha$ decreases below the transition value, $\alpha_{\rm T}$, the solution through the outermost critical point takes-over. At $\alpha=\alpha_{\rm T}$, the two types of solutions merge.
The transition value $\alpha_{\rm T}$ depends on the value of $\beta$, and can be derived by imposing the condition that the trans-sonic flow passes through both the innermost and the outermost critical point. The transition $\alpha$ then is,
@@
where $\eta=(1+1/c)\ln(1+c)-1$. The transition $\alpha$ is shown as a dashed curve in Fig. 5. For a seed black hole of $10$ M, and a halo mass $10^{12}$ M, the transition $\alpha$ is $\approx 0.02$. The transition from one solution to the other leads to two distinct phases of black hole growth, as we discuss in the next section.
@@ 3.1 The black hole mass accretion rate in an NFW halo
By substituting $x_{c}=\beta/2$ for the innermost critical point, and, $x_{c}\approx(1+\beta)/2$ for the outermost critical point which is outside the virial radius, we retrieve the following expression revealing two phases of mass accretion,
@@
The dimensionless mass accretion rate, $\lambda$, is plotted in Fig. 4 and Fig. 5 in the parameter space of $\alpha$ and $\beta$. The initial phase (for $\alpha\gg\alpha_{\rm T}$) exhibits a low mass accretion rate and the black hole mass remains almost the same as the initial seed mass. The initial ‘slow’ phase of growth is set by the solution through the innermost critical point that is due to a high $\alpha$ (Fig. 3 bottom panel). As the value of $\alpha=c_{s}^{2}R_{s}/(GM_{\rm h})$ decreases with redshift due to increasing halo mass, at a specific $\alpha$ (Eq. 24 and the dashed line in Fig. 5) the solution transitions to the one through the outermost critical point (Fig. 3 top panel). Then the second phase of a ‘rapid’ growth with a high mass accretion rate begins (flat line in Fig. 4 and the red zone in Fig. 5). In this phase, the dimensionless mass accretion rate, $\lambda$, is constant, and the actual accretion rate is proportional to $M_{\rm h}^{2}$.
The phases of evolution are also proposed heuristically by Li (2012); and with cosmological simulations (e.g. Bower et al., 2017; Dubois et al., 2015; Anglés-Alcázar et al., 2017; McAlpine et al., 2018) however the origin of phases reported by the simulations is different from ours due to two reasons. Firstly the seed black hole mass used in the simulations is $\sim 10^{5}$ M (Dubois et al., 2015; Bower et al., 2017; McAlpine et al., 2018), while in our work the black hole grows from an initial mass of $10$ M, typical for a stellar black hole. Secondly, the simulations tie the phases of SMBH growth with supernovae feedback (Dubois et al., 2015; Anglés-Alcázar et al., 2017). Bower et al. (2017) attributes the initial slow phase to the clearance of gas by ‘buoyant’ outflows, and the rapid ‘nonlinear’ phase to the failure of outflows due to increased virial temperature of the halo that results in availability of gas in the halo (see also McAlpine et al., 2017, 2018). However, in our work the origin of the phases of black hole growth are purely hydrodynamic.
Figure 5: The dimensionless mass accretion rate onto a black hole at the centre of a NFW halo, $\log_{10}\lambda$ (Eq. 26) shown in colour as a function of $\alpha=R_{s}/l_{h}$ and $\beta=M_{\rm bh}/M_{\rm h}$, where $R_{s}$ is the NFW scale radius related to the virial radius through a concentration parameter $c$ as $R_{s}=R_{\rm vir}/c$, and we take $c=5$. The mass accretion rate is low and it is proportional to $\beta^{2}$ for high $\alpha$, and it is constant ($\approx 1.12$) for low values of $\alpha$. The two regimes (phases) are separated by a transition $\alpha$ (Eq. 24) shown as a dashed curve.
@@ 4 Growth of the central black hole
In this section we study the growth of a black hole in high redshift haloes, as a consequence of the accretion rate derived in the previous section.
Accretion rate depends on the temperature and the density of the gas available for accretion far away. It also depends on the mass of the dark matter halo, that evolves with redshift as, $M_{\rm h}(z)=M_{\rm h,0}(1+z)^{a}e^{-bz}$, where $a$ and $b$ depend on $M_{\rm h,0}$. With this, Eq. 25 can be written in the following differential equation form,
@@
where, $\rho_{\infty,0}=\rho_{\infty}(1+z)^{-3}$. We use $dt/dz=-(1+z)^{-1}H(z)^{-1}$, and the accretion histories from Correa et al. (2015a) to solve the above differential equation, with the latest cosmological parameters from Planck Collaboration et al. (2020). Figure 6 (top panel, solid black curve) shows the evolution of a seed black hole of $10$ M in a massive halo $M_{\rm h,0}=10^{13}$ M, for a flow temperature $T=10^{4}$ K and a fiducial value of density far away, $\rho_{\infty}=2\times 10^{-3}\rho_{\rm b}$, where $\rho_{\rm b}$ is the cosmic baryon density. The shaded region shows the range when varying the density from $0.1\rho_{\infty}$ to $10\rho_{\infty}$. The two phases of accretion are evident (black curve), as the black hole has negligible growth upto a redshift of $\approx 12.5$, but thereafter its mass increases rapidly to supermassive scales. At redshift $8$ it grows to $\approx 10^{8}$ M. The predicted evolution as shown by the black curve and the shaded region agrees with the masses of the SMBHs detected recently by the JWST (Larson et al., 2023; Maiolino et al., 2023a, b; Kokorev et al., 2023) and from earlier detection of quasars (Bañados et al., 2016, 2018; Mortlock et al., 2011; Reed et al., 2019; Matsuoka et al., 2018b, a, 2019b, 2019a; Yang et al., 2020; Wang et al., 2021; Inayoshi et al., 2020).
@@
Figure 6: Top panel: mass of the growing black hole, $M_{\rm bh}$ with redshift, $z$. The results are shown for a seed black hole of mass $10$ M and for an accretion flow with $T=10^{4}$ K and $\rho_{\infty}=2\times 10^{-3}\rho_{\rm b}$ where $\rho_{\rm b}$ is the cosmic baryon density. The shaded region shows the range when varying $\rho_{\infty}$ by multiplication factor of $0.1-10$. The magenta circles are for supermassive black holes (SMBHs) detected recently by the JWST (Larson et al., 2023; Maiolino et al., 2023a, b; Kokorev et al., 2023); the green square is for the SMBH in a quasar detected by ALMA (Wang et al., 2021); black triangles are for the SMBHs detected by WISE (Bañados et al., 2018; Yang et al., 2020); red down-pointing triangle is for the SMBH mass determined from quasar observation using UKIRT (Mortlock et al., 2011), blue diamonds are for the detection by VHS (Reed et al., 2019). The red small plus symbols are for the quasars detected at $z\geq 6$ complied by Bañados et al. (2016), and orange small cross symbols are for the quasars from Matsuoka et al. (2018b, a, 2019b, 2019a); see also Inayoshi et al. 2020 for a compilation of the quasar data. Bottom panel: Eddington ratio corresponding to the black curve in the top panel calculated for an radiative efficiency, $\varepsilon=0.05$, compared with the measured values from recent observations shown as datapoints with errorbars (symbols for references are the same as in the top panel). The shaded region shows the possible range if $\varepsilon$ is varied from $0.01$ to $0.1$.
@@
We also calculate the Eddington ratio for the black holes in this work (Fig. 6, bottom panel). For that we adopt the standard relation, $L=\varepsilon\dot{M}_{\rm bh}c^{2}$, to estimate the luminosity for an SMBH. Then the Eddington ratio is simply given by, $\Gamma=L/L_{\rm Edd}$, where $L_{\rm Edd}=[1.3\times 10^{38}~{}{\rm erg~{}s^{-1}}]$ $M_{\rm bh}/{\rm M}_{\odot}$.
@@
Figure 6 (bottom panel) shows the evolution of the Eddington ratio, $\Gamma$, as a solid black curve corresponding to the growth of a seed black hole of $10$ M shown with a black curve in the top panel. We use a radiative efficiency of $\varepsilon=0.05$, and also show the effects of its variation from $0.01$ to $0.1$ (Inayoshi et al., 2020, and references therein). The Eddington ratio, $\Gamma$, is low to begin with at high redshift ($z>13$) and then it increases to super-Eddington values for a short period of roughly $100$ Myr. This period is likely to be even shorter as the radiative efficiency may be lower due to photon trapping (Begelman, 1979) or magnetic-field effects (Narayan et al., 2003; Sądowski et al., 2015; McKinney et al., 2015).
@@
The accretion flow may not continue if there is a strong radiative feedback during the super-Eddington episode. However, the super-Eddington episode is short-lived (roughly 100 Myr), and more importantly the black hole mass is low during that episode, which may not be adequate to drive outflows and stop accretion (Silk & Rees, 1998; King, 2003; Li, 2012). The minimum mass required to have significant radiative feedback (King, 2003) is shown as a red dashed line in the top panel of Fig. 6. The Eddington ratio is briefly greater than one (bottom panel), but then the solid curve for the black hole growth lies below the dashed line (top panel). We conclude that the black hole growth is fine-tuned in such a way that the short episode of super-Eddington accretion coincides with a low black hole mass that is inadequate to curtail the accretion.
@@
The gas infall into a halo is an essential ingredient for the formation of the galaxy as it builds a gas reservoir from which stars form (e.g. Lilly et al., 2013; Schaye et al., 2015; Lacey et al., 2016; Vogelsberger et al., 2020; Sharma & Theuns, 2020; Trapp et al., 2022). It is debatable when the infalling gas is primarily used for the growth of the black hole or when it is used for the formation of stars and the galaxy (e.g. Di Matteo et al., 2008; Volonteri et al., 2015; Mo et al., 2023; Mountrichas, 2023). Our solution shows that the initial accretion rate onto the black hole is negligible; therefore, in that period, most of the cosmological infall is likely utilized for the formation of the galaxy. However, during the latter rapid accretion phase, the accretion rate onto the black hole is very high, and then most of the cosmological infall is likely consumed by the black hole.
@@
The growth of the black holes in this paper should not be extrapolated for redshifts, $z\leq 6$, since that would indicate unusually high black hole masses ($\geq 10^{9}$ M). It is well known that the giant black holes with Eddington ratios close to unity drive outflows, and as a result, their growth is limited (e.g. Churazov et al., 2005; King, 2010; Bischetti et al., 2022; Capelo et al., 2023). The AGN radiation-driven outflows are a remedy to explain the mismatch between the halo mass function and the galaxy luminosity function at the high mass end (Benson et al., 2003; Bower et al., 2006; Croton et al., 2006; Wechsler & Tinker, 2018). The quasars, including the ones recently detected by the JWST, exhibit signatures of outflows (Bischetti et al., 2022; Cresci et al., 2023; Veilleux et al., 2023). The outflows can inhibit the cosmological infall and hamper the star formation as well (e.g. Bower et al., 2017; Costa et al., 2018). For the highest mass SMBHs, a delicate balance likely develops between the growth of the SMBH, halo, and the galaxy (e.g. Sharma & Theuns, 2020) that leads to a secular co-evolutionary behaviour (e.g. Gebhardt et al., 2000; Ferrarese, 2002; Heckman & Best, 2014; Li et al., 2022).
5 Discussions and Conclusions

### v3

@@ 3 Accretion onto a seed black hole in an NFW dark matter halo
In a Hernquist halo, the mass remains finite when the density profile is integrated up to infinity. However, for the NFW halo, the mass diverges when integrating the density profile to infinity. Therefore, the halo’s mass is defined up to a virial radius at which the halo profile is truncated. The virial radius is defined as (e.g. Mo et al., 1998),
@@
Locations of the critical points for specific values of $\alpha$ and $\beta$ are shown in Fig. 2 bottom panel. There are three critical points: innermost, intermediate, and outermost.
@@
The outermost critical point is difficult to retrieve analytically. However, we can use the fact that it is approximately at the virial radius or outside it (Fig. 2, see also Hobbs et al. 2012), and the NFW halo is truncated at the virial radius. Therefore, the first term on the right-hand side of Eq. 23 is $1/x_{c}^{2}$, which yields $x_{c}\approx(1+\beta)/2$, same as in the case of the Hernquist halo.
In Fig. 3, we show a comparison of the trans-sonic solutions for a Hernquist halo and an NFW halo (for $c=5$). The two types of trans-sonic solutions are shown: the solution through the innermost critical point (bottom panel) and through the outermost critical point (top panel). Solutions through the innermost point for the two types of dark matter haloes are almost identical because the innermost point is due to the black hole. Moreover, the solutions through the outermost critical point differ only slightly between the two haloes.
@@
For both the Hernquist and the NFW halo, there are three critical points: innermost, intermediate, and outermost (Fig. 2). The Mach number, $\mathcal{M}$, for the accretion flow through the outermost critical point (top panel of Fig. 3), stops increasing as $x$ is closer to the location of the intermediate critical point. Then, there are three possibilities depending on the value of $\alpha$. First, for very low $\alpha$, $\mathcal{M}$ increases again as $x$ approaches the location of the innermost critical point. Second, when $\alpha$ is higher and equal to a transition value ($\alpha_{\rm T}$), $\mathcal{M}$ decreases and passes through the innermost critical point. Third, for very high $\alpha$, $\mathcal{M}$ decreases rapidly, and the solution cannot reach the innermost critical point; instead, it loops around the intermediate critical point.
In the third case, the solution through the outermost critical point cannot reach the black hole. However, the accretion occurs as the trans-sonic solution through the innermost critical point takes over (Fig. 3 bottom panel).
The value of $\alpha$ decreases with redshift, and we can also understand the switch-over between the two solutions from the perspective of a decreasing $\alpha$. For a very high value of $\alpha$, the feasible trans-sonic solution for the growth of the black hole is through the innermost critical point. As $\alpha$ decreases below the transition value, $\alpha_{\rm T}$, the solution through the outermost critical point takes over. At $\alpha=\alpha_{\rm T}$, the two types of solutions merge.
The transition value $\alpha_{\rm T}$ depends on the value of $\beta$ and can be derived by imposing the condition that the trans-sonic flow passes through both the innermost and the outermost critical point. The transition $\alpha$ then is,
@@
where $\eta=(1+1/c)\ln(1+c)-1$. The transition $\alpha$ is shown as a dashed curve in Fig. 5. For a seed black hole of $10$ ${}_{\odot}$, and a halo mass $10^{12}$ ${}_{\odot}$, the transition $\alpha$ is $\approx 0.02$. The transition from one solution to the other leads to two distinct phases of black hole growth, as we discuss in the next section.
@@ 3.1 The black hole mass accretion rate in an NFW halo
By substituting $x_{c}=\beta/2$ for the innermost critical point and $x_{c}\approx(1+\beta)/2$ for the outermost critical point that is outside the virial radius, we retrieve the following expression revealing two phases of mass accretion:
@@
The dimensionless mass accretion rate, $\lambda$, is plotted in Fig. 4 and Fig. 5 in the parameter space of $\alpha$ and $\beta$. The initial phase (for $\alpha\gg\alpha_{\rm T}$) exhibits a low mass accretion rate, and the black hole mass remains almost the same as the initial seed mass. The initial ‘slow’ phase of growth is set by the solution through the innermost critical point that is due to a high $\alpha$ (Fig. 3 bottom panel). As the value of $\alpha=c_{s}^{2}R_{s}/(GM_{\rm h})$ decreases with redshift due to increasing halo mass, at a specific $\alpha$ (Eq. 24 and the dashed line in Fig. 5) the solution transitions to the one through the outermost critical point (Fig. 3 top panel). Then, the second phase of a ‘rapid’ growth with a high mass accretion rate begins (flat line in Fig. 4 and the red zone in Fig. 5). In this phase, the dimensionless mass accretion rate, $\lambda$, is constant, and the actual accretion rate is proportional to $M_{\rm h}^{2}$.
The phases of evolution are also proposed heuristically by Li (2012); and with cosmological simulations (e.g. Bower et al., 2017; Dubois et al., 2015; Anglés-Alcázar et al., 2017; McAlpine et al., 2018); however, the origin of phases reported by the simulations is different from ours due to two reasons. Firstly, the seed black hole mass used in the simulations is $\sim 10^{5}$ ${}_{\odot}$ (Dubois et al., 2015; Bower et al., 2017; McAlpine et al., 2018), while in our work, the black hole grows from an initial mass of $10$ ${}_{\odot}$, typical for a stellar black hole. Secondly, the simulations tie the phases of SMBH growth with supernovae feedback (Dubois et al., 2015; Anglés-Alcázar et al., 2017). Bower et al. (2017) attributes the initial slow phase to the clearance of gas by ‘buoyant’ outflows and the rapid ‘nonlinear’ phase to the failure of outflows due to increased virial temperature of the halo that results in the availability of gas in the halo (see also McAlpine et al., 2017, 2018). However, in our work, the origin of the phases of black hole growth is purely hydrodynamic.
Figure 5: The dimensionless mass accretion rate onto a black hole at the centre of an NFW halo, $\log_{10}\lambda$ (Eq. 26) shown in colour as a function of $\alpha=R_{s}/l_{h}$ and $\beta=M_{\rm bh}/M_{\rm h}$, where $R_{s}$ is the NFW scale radius related to the virial radius through a concentration parameter $c$ as $R_{s}=R_{\rm vir}/c$, and we take $c=5$. The mass accretion rate is low, proportional to $\beta^{2}$ for high $\alpha$, and it is constant ($\approx 1.12$) for low values of $\alpha$. The two regimes (phases) are separated by a transition $\alpha$ (Eq. 24) shown as a dashed curve.
@@ 4 Growth of the central black hole
In this section, we study the growth of a black hole in high redshift haloes as a consequence of the accretion rate derived in the previous section.
The accretion rate depends on the temperature and the density of the gas available for accretion far away. It also depends on the mass of the dark matter halo that evolves with redshift as, $M_{\rm h}(z)=M_{\rm h,0}(1+z)^{a}e^{-bz}$, where $a$ and $b$ depend on $M_{\rm h,0}$. With this, Eq. 25 can be written in the following differential equation form,
@@
where, $\rho_{\infty,0}=\rho_{\infty}(1+z)^{-3}$. We use $dt/dz=-(1+z)^{-1}H(z)^{-1}$, and the accretion histories from Correa et al. (2015a) to solve the above differential equation, with the latest cosmological parameters from Planck Collaboration et al. (2020). Figure 6 (top panel, solid black curve) shows the evolution of a seed black hole of $10$ ${}_{\odot}$ in a massive halo $M_{\rm h,0}=10^{13}$ ${}_{\odot}$, for a flow temperature $T=10^{4}$ K and a fiducial value of density far away, $\rho_{\infty}=2\times 10^{-3}\rho_{\rm b}$, where $\rho_{\rm b}$ is the cosmic baryon density. The shaded region shows the range when varying the density from $0.1\rho_{\infty}$ to $10\rho_{\infty}$. The two phases of accretion are evident (black curve), as the black hole has negligible growth up to a redshift of $\approx 12.5$, but after that, its mass increases rapidly to supermassive scales. At redshift $8$, it grows to $\approx 10^{8}$ ${}_{\odot}$. The predicted evolution as shown by the black curve and the shaded region, agrees with the masses of the SMBHs detected recently by the JWST (Larson et al., 2023; Maiolino et al., 2023a, b; Kokorev et al., 2023) and from earlier detection of quasars (Bañados et al., 2016, 2018; Mortlock et al., 2011; Reed et al., 2019; Matsuoka et al., 2018b, a, 2019b, 2019a; Yang et al., 2020; Wang et al., 2021; Inayoshi et al., 2020).
@@
Figure 6: Top panel: mass of the growing black hole, $M_{\rm bh}$ with redshift, $z$. The results are shown for a seed black hole of mass $10$ ${}_{\odot}$ and for an accretion flow with $T=10^{4}$ K and $\rho_{\infty}=2\times 10^{-3}\rho_{\rm b}$ where $\rho_{\rm b}$ is the cosmic baryon density. The shaded region shows the range when varying $\rho_{\infty}$ by a multiplication factor of $0.1-10$. The magenta circles are for supermassive black holes (SMBHs) detected recently by the JWST (Larson et al., 2023; Maiolino et al., 2023a, b; Kokorev et al., 2023); the green square is for the SMBH in a quasar detected by ALMA (Wang et al., 2021); black triangles are for the SMBHs detected by WISE (Bañados et al., 2018; Yang et al., 2020); red down-pointing triangle is for the SMBH mass determined from quasar observation using UKIRT (Mortlock et al., 2011), blue diamonds are for the detection by VHS (Reed et al., 2019). The red small plus symbols are for the quasars detected at $z\geq 6$ compiled by Bañados et al. (2016), and orange small cross symbols are for the quasars from Matsuoka et al. (2018b, a, 2019b, 2019a); see also Inayoshi et al. 2020 for a compilation of the quasar data. Bottom panel: Eddington ratio corresponding to the black curve in the top panel calculated for a radiative efficiency, $\varepsilon=0.05$, compared with the measured values from recent observations shown as datapoints with errorbars (symbols for references are the same as in the top panel). The shaded region shows the possible range if $\varepsilon$ is varied from $0.01$ to $0.1$.
@@
We also calculate the Eddington ratio for the black holes in this work (Fig. 6, bottom panel). For that, we adopt the standard relation, $L=\varepsilon\dot{M}_{\rm bh}c^{2}$, to estimate the luminosity for an SMBH. Then the Eddington ratio is simply given by, $\Gamma=L/L_{\rm Edd}$, where $L_{\rm Edd}=[1.3\times 10^{38}~{}{\rm erg~{}s^{-1}}]$ $M_{\rm bh}/{\rm M}_{\odot}$.
@@
Figure 6 (bottom panel) shows the evolution of the Eddington ratio, $\Gamma$, as a solid black curve corresponding to the growth of a seed black hole of $10$ ${}_{\odot}$ shown with a black curve in the top panel. We use a radiative efficiency of $\varepsilon=0.05$, and also show the effects of its variation from $0.01$ to $0.1$ (Inayoshi et al., 2020, and references therein). The Eddington ratio, $\Gamma$, is low at the beginning at high redshift ($z>13$), and then it increases to super-Eddington values for a short period of roughly $100$ Myr. This period is likely even shorter as the radiative efficiency may be lower due to photon trapping (Begelman, 1979) or magnetic-field effects (Narayan et al., 2003; Sądowski et al., 2015; McKinney et al., 2015).
@@
The accretion flow may not continue if there is strong radiative feedback during the super-Eddington episode. However, the super-Eddington episode is short-lived (roughly 100 Myr), and more importantly, the black hole mass is low during that episode, which may not be adequate to drive outflows and stop accretion (Silk & Rees, 1998; King, 2003; Li, 2012). The minimum mass required to have significant radiative feedback (King, 2003) is shown as a red dashed line in the top panel of Fig. 6. The Eddington ratio is briefly greater than one (bottom panel), but then the solid curve for the black hole growth lies below the dashed line (top panel). We conclude that the black hole growth is fine-tuned in such a way that the short episode of super-Eddington accretion coincides with a low black hole mass that is inadequate to curtail the accretion.
@@
The gas infall into a halo is an essential ingredient for the formation of the galaxy as it builds a gas reservoir from which stars form (e.g. Lilly et al., 2013; Schaye et al., 2015; Lacey et al., 2016; Vogelsberger et al., 2020; Sharma & Theuns, 2020; Trapp et al., 2022). It is debatable when the infalling gas is primarily used for the growth of the black hole or when it is used for the formation of stars and the galaxy (e.g. Di Matteo et al., 2008; Volonteri et al., 2015; Mo et al., 2023; Mountrichas, 2023). Our solution shows that, in the initial phase, the accretion rate onto the black hole is negligible; therefore, in that phase, most of the cosmological infall is likely utilised to form the galaxy. However, during the latter rapid accretion phase, the accretion rate onto the black hole is very high, and then most of the cosmological infall is likely consumed by the black hole.
@@
The growth of the black holes in this paper should not be extrapolated for redshifts, $z\leq 6$, since that would indicate unusually high black hole masses ($\geq 10^{9}$ ${}_{\odot}$). It is well known that the giant black holes with Eddington ratios close to unity drive outflows, and as a result, their growth is limited (e.g. Churazov et al., 2005; King, 2010; Bischetti et al., 2022; Capelo et al., 2023). The AGN radiation-driven outflows are a remedy to explain the mismatch between the halo mass function and the galaxy luminosity function at the high mass end (Benson et al., 2003; Bower et al., 2006; Croton et al., 2006; Wechsler & Tinker, 2018). The quasars, including the ones recently detected by the JWST, exhibit signatures of outflows (Bischetti et al., 2022; Cresci et al., 2023; Veilleux et al., 2023). The outflows can inhibit the cosmological infall and hamper the star formation as well (e.g. Bower et al., 2017; Costa et al., 2018). For the highest mass SMBHs, a delicate balance likely develops between the growth of the SMBH, halo, and the galaxy (e.g. Sharma & Theuns, 2020) that leads to a secular co-evolutionary behaviour (e.g. Gebhardt et al., 2000; Ferrarese, 2002; Heckman & Best, 2014; Li et al., 2022).
5 Summary and Conclusions